the peak of the mountain is 9,000 meters from the base. To climb up the side of the mountain it is 12,000 meters. how many meters deep is the radius of the mountain at its base?
1
Expert's answer
2016-02-04T04:06:50-0500
If the peak of the mountain is 9,000 meters from the base and the side of the mountain it is 12,000 meters, then to find the radius of the mountain at its base we need to calculate it as one of the sides of the rectangular triangle.
Base^2 = Side^2 - Height^2 = 12,000^2 - 9,000^2 = 63,000,000,
Base = √63,000,000 = 7,937.25 meters.